OpenStudy (anonymous):

A mole of hydrogen atoms with their electron in the n = 6 energy level emit 109.4 kJ of energy. In what energy level (n) does the electron end up?

OpenStudy (cuanchi):

\[Delta E=-2.18 \times 10^{-18}J(\frac{ 1 }{ n _{f}^{2} }-\frac{ 1 }{ n _{i}^{2} })\]

OpenStudy (anonymous):

thanks for the formula but what does n^2 f mean

OpenStudy (cuanchi):

that is the value that you have to calculate is the final level where the electron felt.

OpenStudy (anonymous):

oh okay so i know my energy and my final but am solving for the inital?

OpenStudy (anonymous):

or is n=6 my initial.. I'm not sure?

OpenStudy (cuanchi):

yes n=6 is your initial and you have to find your final. But the formula is for the energy of a single electron, and in your problem you have a mol of electron so you have to divide that energy by the Avogadro's number

OpenStudy (cuanchi):

Besides that energy is in kJ and in the formula the constant is in Joules